OpenStudy (anonymous):

if f(x)=x and g(x)= 2x-1, find (f-g)(x). (1 point) --- x-3 a) -2x^2+8x-3 ------------ x-3 b) -2x^2+6x+1 ------------ x-3 c) -2x^2+5x-3 ------------ x-3 d) 2x^2+6x+3 ------------ x-3

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

@CoolKatt16

OpenStudy (anonymous):

dont know sorry :P

hero (hero):

\[(f - g)(x) = f(x) - g(x) = \frac{x}{x - 3} - (2x - 1)\]

OpenStudy (anonymous):

@Hero so is it a,b,c or d??

hero (hero):

You have to perform the subtraction completely to figure it out.

OpenStudy (anonymous):

@Hero i don't know how

hero (hero):

Suppose you had to subtract \(\dfrac{24}{5} - 4\). How would you do it?

OpenStudy (anonymous):

@Hero make it 24/5 - 4/1?

hero (hero):

And then what would you do?

OpenStudy (anonymous):

@Hero it would be 20/4?

hero (hero):

Hmmm. So basically, what you are saying is you don't know how to subtract regular fractions. You should review how to subtract basic fractions before dealing with algebraic fractions.